40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Nov 18 15:45:11 2011, Mirko, Update, IOO, Mode cleaner noise projection WholeMCNoiseProjection.pngWholeMCNoiseProjection.fig
    Reply  Fri Nov 18 19:57:19 2011, Mirko, Update, IOO, Mode cleaner noise projection MC_ELP.pngWithAndWithoutHPs.pdfSeismCoh.pdf
       Reply  Fri Nov 18 23:44:33 2011, rana, Update, IOO, Mode cleaner noise projection 
          Reply  Sat Nov 19 01:26:16 2011, Den, Update, IOO, Mode cleaner noise projection 
          Reply  Sat Nov 19 12:57:55 2011, Mirko, Update, IOO, Mode cleaner noise projection CouplingMClengthsToMCF.pdfNpModeCleaner.pdf
Message ID: 5952     Entry time: Fri Nov 18 19:57:19 2011     In reply to: 5949     Reply to this: 5953
Author: Mirko 
Type: Update 
Category: IOO 
Subject: Mode cleaner noise projection 

 

Some more info on this:
 

f > 1 Hz:

At these frequencies the pendulum should be quieter than the stacks. By quite a bit actually since there is the stack resonance at a couple Hz. 'Glueing' them together via the local control is not wise. We put an elliptic LP ( 2.5Hz, 4th order 6dB) into the C1:SUS-MC?_SUSPOS pathes and MC-F got better above 1Hz

MC_ELP.png

Added an extra LP @ 10 Hz afterwards. Doesn't make a visible difference.

f < 1 Hz:

Now here is more stuff to consider.

1. The OSEMs glue the MC mirrors to the stacks
2. The pendulum TF should be 1
3. It shouldn't matter if the OSEMs do or do not act on the mirror at these frequencies, assuming they don't add extra noise.
4. Page http://nodus.ligo.caltech.edu:8080/40m/5547 seems to indicate OSEM sensor noise is so low it can be neglected.

Reduced OSEM gain below 1Hz:

If we reduce the gain in the OSEMs by adding additional HP filters ( cheby2, HP 0.3Hz, 6dB 4th order ) the happens:

1. MC length gets a bit more noisy at low frequencies - should be looked into some more
2. Coherence between the GUR1 seismometer and MC length goes up between 1E-2 and 1E-1 Hz:

( Ref is with low OSEM gain )

WithAndWithoutHPs.pdf

Possible explanation:

The stacks might be more correlatedly moving together than the pendulums. This would be not so nice for OAF test, but really fine for actually using the MC.
Todo: Measure the OSEM to seismometer coherences with high and low OSEM gains.

For reference the seismometer coherence with one another:
SeismCoh.pdf

 

 

ELOG V3.1.3-